PIANOS conclusion document

Group Linja

Helsinki 31st August 2005 Software Engineering Project UNIVERSITY OF HELSINKI Department of Computer Science Course

581260 Software Engineering Project (6 cr)

Project Group Joonas Kukkonen Marja Hassinen Eemil Lagerspetz

Client Marko Salmenkivi

Project Masters Juha Taina Vesa Vainio (Instructor)

Homepage

http://www.cs.helsinki.fi/group/linja

Contents

1 Introduction 1 1.1 Version history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Schedule review 2 2.1 Milestones . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

3 Working hour reports 3

4 Products 4 4.1 Documents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4.2 Executable prototypes . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4.3 Final program source code . . . . . . . . . . . . . . . . . . . . . . . . . 4

5 Risk review 5

6 Lessons learned during the project 6 6.1 Eemil Lagerspetz . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 6.2 Marja Hassinen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 6.3 Joonas Kukkonen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

7 References 8

Appendices

1 Working hour lists

1 Introduction

This is a conclusion document that details the products and other results of the PIANOS project. This document effectively ends the project. In this document the project schedule is outlined and compared to the actual realized schedule. The working hours used for the project are summarized and a final report of realized risks is compared to the risk assessment done in the early phases of the project. Finally, project members write a little on lessons learned during the project.

1.1 Version history

Version Date  Modifications
0.2  26.08.2005 Document template
1.0  30.08.2005 First full-featured draft
1.1  31.08.2005 Reviewed and corrected final

2 Schedule review

2.1 Milestones

This chapter represents the estimated dates and the correct dates of the products (documents and the software).

Deliverable  Estimated date Correct date
Project plan (final)  27.5.  27.5.
Software requirements specification  17.6.  21.6.
Software design document  22.7.  3.8.
Software product (executable version)  19.8.  around 16.8.
Software product (a version working correctly) -  28.8.
Implementation document  -  30.8.
Manual  -  25.8.
Testing document  -  30.8.
Final software product (release)  2.9  2.9.

3 Working hour reports

Statistics

Week Group members Total working hours Avg. hours per person
1  5  65  13
2  5  79  16
3  4  56  14
4  4  73  18
5  4  65  16
6  4  53  13
7  4  56  14
8  4  82  21
9  3  80  27
10  3  62  21
11  3  60  20
12  3  66  22
13  3  75  25
14  3  ?  ?
Total  ?  ?

The working hour listings are included as an appendix.

4 Products

This section describes the products of the project and their sizes.

4.1 Documents

Document name  Pages
Project plan  25
Requirements specification 36
Design document  70
User manual  34
Implementation document 49
Testing document  76
Conclusion document  ?

4.2 Executable prototypes

Model name  Lines of code
The simple bird model  1049
Spatial bird model with spatial dependence 1003

Before these, two other prototypes were made but they had no semantic background and therefore didn’t produce any results.

4.3 Final program source code

Package name  Lines of code Executable lines of code
PIANOS.exceptions  67  12
PIANOS  315  169
PIANOS.datastructures 1728  612
PIANOS.io  2192  1074
PIANOS.generator  3572  1963

Total lines of code: 7874 Total executable lines of code: 3831 (not including: comment lines, empty lines, curly bracket lines)

5 Risk review

This section presents the risks described in Project plan and their final status, that is, whether the risk has realized or not.

The fields of the following table:

• Risk name: a short description of the risk also found in the project plan

• Probability: the estimated likelihood of the risk coming true, on a scale of ‘Low, Moderate, High’.

• Seriousness: an estimate of impact on the project, ‘High’ being the most serious, ‘Moderate’ and ‘Low’ the other choices.

• Final status: this is the risk’s status at the end of the project

Risk name  Probability Seriousness Final status
Temporary loss of work force  High  Low  Realized
Permanent loss of work force  Low  High  Realized - twice
CS network failure  Moderate Moderate Realized
Loss of project data  Moderate Moderate Not realized
A part of the project is delayed Moderate High Realized
Schedule pressure  Moderate Moderate Realized
Evolution of requirements Low High Not realized
Over-planning  Moderate Moderate Not realized
Over-prototyping  Moderate Moderate Not realized
Harder than expected  Moderate Moderate Not realized
The problem is more extensive than expected Moderate High  Not realized
Incompatibility  Moderate High  Not realized
Fortran issues  Moderate High  Realized to some extent
Mis-implementation  Moderate High  Cannot be determined
Wrong approach  Moderate High  Cannot be determined
Lack of technical specification  Moderate Moderate Not realized

There were no realized risks that were not found out at the project planning phase.

6 Lessons learned during the project

This section contains the experiences of the project members on the project as they have written them.

6.1 Eemil Lagerspetz

In my eyes, the significance of planning has increased as a result of this project. I used to plan only the big ideas very quickly and then program as I thought was appropriate, but in this project we designed in detail and then the implementation could follow the design closely, only a few places needed modifications. I was also suprised that the use of CVS, L^{A}TEXand development environments can make projects like this so much easier. After using CVS I cannot think that the project could have beed done without a versioning management system.

6.2 Marja Hassinen

Project phases

After the Software Engineering course I had a vague picture of project phases but during this project the picture got clearer. Now I know not only what the purpose of each phase is but also the activities included in each phase.

Documents: what to write

At the beginning of the cource one of my goals was to learn to write documents and to know, what kind of things the documents should include. During the course I got a clearer view of the documents that should be written during a software project.

Design: what it is

Before this project I had programmed only small applications and it hadn’t been necessary to design them. At the Software Engineering course there were only a couple of illustration techiniques presented, but I didn’t got a very explicit idea of what design really is. During this project I learned what kind of activities can be done at the design phase and how the ideas about the program get clearer step by step and can be represented to other people.

The problem domain

• Bayesian models: Before this course I knew the basics of Bayesian models but during the course I learned why they are useful and what kind of things can be represented with them.

• Spatial dependencies: During the course I learned what kind of things can be represented with spatial models and what kind of difficulties arise from spatial dependencies.

• The Metropolis-Hastings algorithm: I have learned how this algorithm works and how to program programs that utilize it.

• The models: I had heard about “the bird model” - now I know what it means.

Fortran

During the project I learned the basics of Fortran. It is quite different from the other programming languages I had used.

Tools

The CVS and LateX tools are handy when working as a team with other people. It was very useful to get to know them.

6.3 Joonas Kukkonen

The greatest lesson for me was how to make a larger programs in a group. The course also gave me insights how important a good software engineering process is. I liked the role of project leader and it taught me, among other things, how to plan and lead meetings.

The software engineering course left the project phases and the documenting a little bit unclear. The project clarified these phases, especially requirement analysis and design. I have learned to write better documents and working on the documents has also improved my writing skills in English. The problem domain has taught me about Bayesian models and modeling of real life phenomenon. It has also let me see what kind of problems the researchers work with.

I have never done any prototypes and so at the start I underestimated the value of making them. The project showed how important and helpful making prototypes can be under right circumstances. During the project I learned to use CVS, Latex and some other helpful programs.

8

7 References

1

Appendix 1. Working hour lists

Marja Hassinen

17.5.2005 INST 3.5 Studying Fortran 17.5.2005 MEET 1.0 Meeting with the project group 17.5.2005 INST 0.5 Studying how to use CVS 17.5.2005 INST 0.5 Writing content to the web page 18.5.2005 INST 2.0 Studying Fortran 18.5.2005 OTHE 0.5 Reading project instructions 19.5.2005 MEET 2.5 Meeting with the project group 19.5.2005 PROJ 0.5 Writing about listing the working hours 19.5.2005 INST 0.5 Studying how to use CVS 19.5.2005 INST 0.5 Studying how to use the report.pl-script 19.5.2005 INST 1.0 Studying how to use LaTeX 19.5.2005 PROJ 1.0 Writing the project plan (Project organization) 20.5.2005 PROJ 1.0 Writing the project plan (Schedule) 20.5.2005 MEET 2.5 Meeting with the client and the project group 23.5.2005 MEET 2.0 Meeting with the project group 23.5.2005 PROJ 1.0 Writing the project plan (Schedule) 23.5.2005 INST 0.5 Studying how to use LaTeX 23.5.2005 REQU 1.0 Reading the IEEE SRS document 24.5.2005 MEET 2.0 Meeting with the client and the project group 24.5.2005 INST 1.0 Studying Fortran 25.5.2005 MEET 2.0 Meeting with the project group 25.5.2005 MEET 0.5 Writing notes of the meating into a file 25.5.2005 PROJ 0.5 Writing the project plan 26.5.2005 INST 2.0 Studying Fortran 27.5.2005 PROJ 1.0 Reading the project plan and marking errors 27.5.2005 MEET 2.0 Meeting with the project group (reviewing the project

plan)

27.5.2005 PROJ 0.5 Correcting grammatical mistakes of the plan (with

Eemil)

29.5.2005 REQU 0.5 Preparing the next meeting 29.5.2005 KNOW 0.5 Reading about Markov random fields 30.5.2005 MEET 2.0 Meeting with the project group and the client 30.5.2005 PROJ 1.5 Reading about the FP method and calculating FP’s 31.5.2005 REQU 2.0 Specifying requirements with the project group 31.5.2005 REQU 1.5 Writing the SRS (data & simulation) 1.6.2005 MEET 2.0 Meeting with the project group 1.6.2005 REQU 1.0 Writing the SRS 1.6.2005 INST 0.5 Studying Fortran 2.6.2005 MEET 2.0 Meeting with the group and the client 2.6.2005 REQU 1.0 Writing the SRS (adding requirements) 3.6.2005 INST 2.0 Studying Fortran 4.6.2005 INST 1.0 Studying Fortran 5.6.2005 REQU 0.5 Writing the SRS (Glossary) 6.6.2005 REQU 2.0 Writing the SRS (Glossary & Requirements) 6.6.2005 MEET 2.0 Meeting with the group and the client 7.6.2005 REQU 2.0 Specifying requirements with the project group 7.6.2005 REQU 2.5 Writing the SRS (Requirements) 8.6.2005 MEET 2.0 Meeting with the project group (requ..) 8.6.2005 REQU 1.5 Writing the SRS (Requ & Gloss) 9.6.2005 MEET 2.0 Meeting with the project group and the client 9.6.2005 REQU 1.0 Writing the SRS (correcting errors) 10.6.2005 MEET 2.0 Meeting with the project group 10.6.2005 REQU 0.5 Correcting errors in SRS 13.6.2005 MEET 2.0 Meeting with the group 13.6.2005 OTHE 0.5 Reporting working hours (statistics, metrics system) 13.6.2005 REQU 1.0 Writing the SRS 13.6.2005 INST 0.5 Studying Fortran 14.6.2005 DESI 0.5 Designing the simulation parameter input form 14.6.2005 OTHE 0.5 Reporting working hours (problems with the system) 14.6.2005 MEET 2.0 Meeting with the project group 14.6.2005 REQU 1.0 Reading the SRS 14.6.2005 DESI 0.5 Reading the IEEE SDD document 15.6.2005 MEET 1.0 Meeting with the project group 15.6.2005 DESI 1.0 Designing the simtech input formats with Eemil 15.6.2005 DESI 0.5 Reading the IEEE SDD document 15.6.2005 DESI 0.5 Designing the simtech input formats 16.6.2005 DESI 1.0 Designing the simtech input formats 16.6.2005 INST 1.5 Studying Fortran

17.6.2005 MEET 2.0 Review (SRS)

17.6.2005 REQU 1.0 Correcting errors in SRS 20.6.2005 DESI 1.0 Reading & designing the simtech input formats 20.6.2005 INST 1.0 Studying Fortran 20.6.2005 MEET 2.0 Meeting with the project group 21.6.2005 DESI 4.0 Designing the prototype (with Eemil) 21.6.2005 DESI 1.0 Designing the prototype 21.6.2005 INST 0.5 Studying Fortran 22.6.2005 INST 1.0 Studying Fortran 22.6.2005 DESI 0.5 Designing the prototype 22.6.2005 MEET 2.0 Meeting with the project group 23.6.2005 DESI 4.0 Designing & coding the prototype (with Eemil) 27.6.2005 DESI 1.0 Designing the prototype (pictures and some minor

changes)

27.6.2005 DESI 1.0 Reading about NAG and finding some routines 27.6.2005 MEET 2.0 Meeting with the project group 27.6.2005 DESI 0.5 Reading notes etc. 28.6.2005 DESI 4.0 Designing & coding the prototype (with Eemil) 29.6.2005 DESI 1.5 Writing the design document 29.6.2005 MEET 2.0 Meeting with the project group 29.6.2005 DESI 1.0 Coding the prototype (with Eemil) 30.6.2005 DESI 2.0 Coding the prototype 1.7.2005 MEET 2.0 Meeting with the project group 18.7.2005 OTHE 0.5 Miscellaneous 18.7.2005 DESI 1.0 Coding & documenting the prototype 18.7.2005 INST 0.5 Studying Fortran 18.7.2005 MEET 2.0 Meeting with the project group 18.7.2005 DESI 1.0 Discussing the protype & designing the generator:

probabilities

19.7.2005 DESI 2.0 Writing the design document (probabilities) at school 19.7.2005 DESI 1.5 Writing the design document (probabilities) at home 19.7.2005 INST 1.0 Reading about NAG and trying to get it work 20.7.2005 MEET 0.5 Meeting with the client 20.7.2005 DESI 1.0 Coding the prototype (with Eemil) 20.7.2005 MEET 2.5 Meeting with the project group 20.7.2005 DESI 1.0 Reading and writing the design document 20.7.2005 DESI 0.5 Discussing the linking 20.7.2005 DESI 0.5 Writing the design document (linking) 21.7.2005 DESI 0.5 Creating prototype definition files (model, simtech

parameters, data)

21.7.2005 DESI 3.0 Coding the prototype & writing the design document,

discussing design (at school)

21.7.2005 DESI 2.0 Correcting errors in prototype (problems with NAG) 22.7.2005 MEET 2.0 Meeting with the project group 22.7.2005 DESI 1.5 Debugging the prototype (it works!) 22.7.2005 DESI 3.0 Writing the design document 25.7.2005 DESI 1.0 Reading the design document 25.7.2005 MEET 1.0 Meeting with the project group 25.7.2005 DESI 2.0 Writing the design document (model desciption lan-

guage) at school

25.7.2005 DESI 1.0 Writing the design document at home 25.7.2005 TEST 1.0 Reading and writing the testing document 26.7.2005 TEST 0.5 Discussing testing. 26.7.2005 DESI 1.0 Discussing design and reading and writing the design

document.

26.7.2005 IMPL 1.0 Discussing implementation and writing Distribu-

tion.java.

26.7.2005 IMPL 0.5 Writing Distribution.java (finished). 26.7.2005 DESI 0.5 Reading the design document 27.7.2005 MEET 1.0 Meeting with the project group 27.7.2005 DESI 1.0 Correcting error in Design document 27.7.2005 IMPL 1.0 Implementing Distribution subclasses 27.7.2005 IMPL 1.5 Implementing Distribution subclasses and Variable

setters

27.7.2005 INST 0.5 Installing Java 1.5 and reading about the new features 28.7.2005 IMPL 2.5 Implementing Distribution subclasses 28.7.2005 TEST 0.5 Trying to unit test Distribution subclasses (problems

with everything)

28.7.2005 TEST 0.5 Unit testing Distribution subclasses (seems ok) 28.7.2005 IMPL 0.5 Discussing implementation and adding features to

DistributionFactory 29.7.2005 MEET 1.0 Meeting with the project group 29.7.2005 IMPL 2.0 Correcting miscellaneuous errors in data structures

and implementing generateAcceptationFormula

29.7.2005 IMPL 2.0 Implementing generateAcceptationFormula 31.7.2005 IMPL 1.5 Implementing generateAcceptationFormula 1.8.2005 IMPL 0.5 Implementing generateUpdateFunctionalCode 1.8.2005 MEET 1.0 Meeting with the project group 1.8.2005 IMPL 1.0 Implementing generateUpdateFunctionalCode 1.8.2005 IMPL 1.5 Implementing generateUpdateFunctionalCode 2.8.2005 DESI 1.5 Reading the design document and discussing the FTR 2.8.2005 IMPL 1.0 Implementing generateUpdateFunctionalCode 2.8.2005 TEST 1.0 Writing a test program for generateAcceptationFor-

mula (problems with other modules)

3.8.2005 IMPL 1.0 Adding features to generateFunctionalNewVal-

uesCode

3.8.2005 MEET 1.0 FTR for design document 3.8.2005 DESI 0.5 Correcting errors in design document 3.8.2005 IMPL 1.0 Discussing implementation and testing 3.8.2005 TEST 1.0 Testing generateAcceptationFormula and correcting

bugs.

4.8.2005 IMPL 1.0 Correcting errors in generateAcceptationFormula 4.8.2005 TEST 1.0 Testing generateAcceptationFormula 4.8.2005 TEST 1.0 Testing generateFunctionalNewValuesCode and cor-

recting errors

4.8.2005 IMPL 0.5 Adding features to generateFunctionalNewVal-

uesCode

4.8.2005 IMPL 0.5 Discussing implementation 4.8.2005 IMPL 0.5 Changing distributions and generator to utilize Ar-

rayList<String>s.

5.8.2005 MEET 1.5 Meeting with the project group 5.8.2005 TEST 0.5 Testing the generateFunctionalNewvaluesCode 5.8.2005 IMPL 1.5 Discussing implementation, adding features to proto-

type, adding new methods.

7.8.2005 IMPL 1.5 Implementing generateProposal 8.8.2005 IMPL 1.0 Implementing generateNewValueCode and gener-

ateAcceptationCode

8.8.2005 MEET 1.5 Meeting with the project group 8.8.2005 IMPL 1.5 Refactoring generateAcceptationFormula and the

methods it uses

8.8.2005 IMPL 1.0 Refactoring generateAcceptationFormula and gener-

ateNewValuesFunctionalCode

9.8.2005 IMPL 1.5 Implementing generateNewValuesFunctionalCode

and the methods it uses

9.8.2005 IMPL 0.5 Correcting errors in generator classes 9.8.2005 TEST 0.5 Generating test cases and testing generateAccepta-

tionFormula (doesn’t work)

9.8.2005 IMPL 1.0 Correcting errors in generateAcceptationFormula 9.8.2005 IMPL 0.5 Implementing generateAcceptationCode 10.8.2005 IMPL 2.0 Correcting miscellaneous programming errors in

Generator (now it can be compiled but not yet run)

10.8.2005 OTHE 2.0 Attending the lecture about testing 11.8.2005 MEET 2.0 Meeting with the project group 11.8.2005 IMPL 2.0 Discussing implementation and implementing Output 12.8.2005 IMPL 2.0 Implementing Output 12.8.2005 IMPL 0.5 Correcting miscellaneous programming errors in

Generator (now it can be run.. wow!)

12.8.2005 TEST 0.5 Testing the generator 15.8.2005 IMPL 1.0 Correcting errors in Output and Proposal. Writing

javadoc comments.

15.8.2005 TEST 1.0 Writing the testing document (test models etc.) and

thinking about testing

15.8.2005 TEST 0.5 Trying to get the produced Fortran code to compile 15.8.2005 MEET 3.0 Meeting with the project group 15.8.2005 TEST 0.5 Hunting a Mysterious Bug 15.8.2005 IMPL 0.5 Correcting the Mysterious Bug 16.8.2005 IMPL 0.5 Refining the generateWriteOutput class 16.8.2005 TEST 0.5 Writing testing document (models) 16.8.2005 TEST 2.5 Testing Acceptation, writing the testing document (at

school)

16.8.2005 TEST 1.0 Testing Acceptation, writing the testing document (at

home)

17.8.2005 TEST 1.0 Testing Acceptation, writing the testing document (at

home)

17.8.2005 TEST 2.0 Testing Acceptation, writing the testing document (at

home)

17.8.2005 TEST 1.0 Discussing miscellaneous errors and features recently

found and correcting them

18.8.2005 MEET 2.0 Meeting with the project group 22.8.2005 OTHE 0.5 Reading e-mail and generating the week report. 23.8.2005 OTHE 1.0 Reading the manual 23.8.2005 TEST 1.0 Reading the source code (Acceptation) and discussing

found bugs and other implementation things

23.8.2005 MEET 3.0 Meeting with the project group 23.8.2005 TEST 0.5 Reading the generated source code for cancer model

and looking for bugs. Didn’t find any.

23.8.2005 IMPL 0.5 Adding features and discussing implementation

(problems with functional parameters - again)

24.8.2005 TEST 0.5 Testing Acceptation with cancer model 24.8.2005 TEST 3.0 Testing Acceptation with model_all_func 24.8.2005 OTPR 1.0 Correcting errors in the manual (problems with cvs

and everything), discussing the CD, viewing its contents

24.8.2005 TEST 0.5 Testing Acceptation and discussing its implementa-

tion

25.8.2005 TEST 1.5 Testing Acceptation with the spatial case and writing

the testing document

25.8.2005 OTPR 0.5 Writing the manual and discussing the implementa-

tion document

25.8.2005 OTPR 0.5 Creating a user_dist.f90 to be delivered to the cus-

tomer, reading documents

25.8.2005 TEST 1.0 Writing the testing document 26.8.2005 MEET 1.0 Meeting with the project group 26.8.2005 IMPL 2.5 Writing the implementation document (requirements

& further development)

26.8.2005 OTPR 1.0 Writing the conclusion document (schedule, risks,

products)

26.8.2005 TEST 0.5 Reading and writing the testing document 26.8.2005 OTPR 0.5 Writing the what have I learned thing 28.8.2005 IMPL 0.5 Writing the implementation document 28.8.2005 OTPR 0.5 Writing the conclusion document and what have I

learned thing

28.8.2005 IMPL 0.5 Discussing implementation (efficiency issues) and

reading the parser source code

28.8.2005 TEST 1.0 System testing with the bird model (the proposal distributions are so messed up that the program simply can’t work)

29.8.2005 TEST 1.0 System testing with the bird model (added debug

prints etc)

29.8.2005 MEET 1.0 Meeting with the project group 29.8.2005 IMPL 1.0 Writing the implementation document 29.8.2005 OTPR 0.5 Writing the conclusion document 30.8.2005 IMPL 0.5 Removing debug prints from the generator and cor-

recting errors in the implementation document

30.8.2005 IMPL 1.0 Reading the implementation document 30.8.2005 TEST 0.5 Reading the testing document and the conclusion doc-

ument

31.8.2005 MEET 1.5 Meeting with the project group 31.8.2005 IMPL 1.0 Correcting errors in the implementation document 31.8.2005 OTPR 1.0 Writing the conclusion document and the testing doc-

ument Eemil Lagerspetz

17.05.2005 MEET 1.0 Official meeting 17.05.2005 INST 3.0 Project homepage, latex-install, CVS-install 17.05.2005 INST 1.0 TortoiseCVS-tutorial, etc 18.05.2005 INST 2.0 CVS and binary files, rewriting the tutorial 19.05.2005 MEET 2.5 Official meeting 19.05.2005 INST 2.0 reorganizing CVS among other things 19.05.2005 PROJ 3.0 Writing about the risks 19.05.2005 INST 1.0 Learning about LaTeX 20.05.2005 MEET 2.5 Official meeting 20.05.2005 PROJ 1.0 Writing about communication and data storage 23.05.2005 MEET 2.0 Official meeting 23.05.2005 PROJ 2.0 Writing parts of the planning document 24.05.2005 MEET 2.0 Official meeting 24.05.2005 PROJ 3.0 Discussing risks with the project manager 25.05.2005 MEET 2.0 Official meeting 25.05.2005 PROJ 1.0 Tweaking the risks 25.05.2005 PROJ 2.0 Adding more content to the risks 25.05.2005 INST 2.0 Configuring a backup CVS server for emergencies 27.05.2005 PROJ 1.0 Reading up for the review of the project plan 27.05.2005 PROJ 1.0 Correcting some errors in the project plan 27.05.2005 PROJ 0.5 Fixing grammatical errors in the project plan 30.05.2005 MEET 2.0 Official meeting 30.05.2005 PROJ 2.0 Writing about testing 31.05.2005 REQU 2.0 Meeting to discuss requirements and use cases 31.05.2005 REQU 1.0 Writing the first draft of the use cases 31.05.2005 REQU 3.0 Discussing and writing the requirements 01.06.2005 MEET 2.0 Official meeting 01.06.2005 REQU 2.0 Working on the use cases and requirements 02.06.2005 MEET 2.0 Official meeting 06.06.2005 MEET 2.0 Official meeting 06.06.2005 REQU 1.0 Adding requirement Stabilities and Ids 07.06.2005 REQU 2.0 Meeting on the srs 07.06.2005 REQU 4.0 Copying and styling diagrams for the srs 08.06.2005 MEET 2.0 Official meeting 08.06.2005 REQU 4.0 Drawing system models, writing glossary 09.06.2005 REQU 2.0 Meeting on the requirements 10.06.2005 MEET 2.0 Official meeting 11.06.2005 REQU 2.0 Redrawing diagrams, writing model description 13.06.2005 MEET 2.0 Official meeting 13.06.2005 REQU 2.0 modifying diagrams, fixing some srs errors 14.06.2005 DESI 2.0 Calculation example meeting 14.06.2005 REQU 1.0 Reading up for the SRS inspection 15.06.2005 MEET 2.0 Official meeting 17.06.2005 REQU 3.0 SRS document inspection 17.06.2005 REQU 2.0 Correcting the SRS 17.06.2005 DESI 2.0 Writing the design document 20.06.2005 MEET 2.0 Official meeting 21.06.2005 DESI 5.5 Designing the prototype, etc. 22.06.2005 MEET 2.0 Official meeting 23.06.2005 DESI 4.0 Designing the prototype 27.06.2005 MEET 2.0 Official meeting 27.06.2005 DESI 1.0 Meeting up with the customer about blocking 27.06.2005 DESI 1.0 Structuring up and modifying the design doc 28.06.2005 DESI 4.0 Designing the prototype, writing the dd 28.06.2005 DESI 6.0 Making the prototype runnable 29.06.2005 MEET 2.0 Official meeting 29.06.2005 DESI 1.0 Redesigning the prototype internals 29.06.2005 DESI 1.5 Fixing prototype bugs 01.07.2005 MEET 2.0 Official meeting 18.07.2005 MEET 2.0 Official meeting 19.07.2005 DESI 2.0 Designing at the department 19.07.2005 DESI 1.0 Designing at home 20.07.2005 DESI 0.5 Meeting on the prototype with the customer 20.07.2005 DESI 1.0 Modifying the prototype for the new model 20.07.2005 DESI 2.5 Design meeting 20.07.2005 INST 1.0 Implementing subsubsubsection 20.07.2005 DESI 1.0 Modifying the prototype 21.07.2005 DESI 3.0 Modifying the prototype 21.07.2005 DESI 0.5 Debugging the prototype 22.07.2005 MEET 2.0 Official meeting 22.07.2005 DESI 1.5 Debugging the prototype 22.07.2005 DESI 3.0 Writing the design document 23.07.2005 DESI 2.0 Writing the design document 25.07.2005 MEET 1.0 Meeting with the project group 25.07.2005 DESI 2.0 Designing the parser 25.07.2005 DESI 3.0 Designing the parser(at home) 26.07.2005 TEST 0.5 Discussing testing 26.07.2005 DESI 1.0 Discussing design, writing the design document 26.07.2005 IMPL 1.0 Discussing implementation and writing the parser 26.07.2005 DESI 2.0 Reading the design document for the review 27.07

27.07.2005 DESI 1.0 The review

27.07.2005 DESI 2.0 Correcting the SDD, writing the parser 27.07.2005 DESI 2.0 Correcting the SDD, writing the parser(at home) 28.07.2005 IMPL 6.0 Writing the parser and other classes 29.07.2005 DESI 1.0 Putting up stuff into the Design document 29.07.2005 IMPL 2.0 Implementing the Parser 29.07.2005 IMPL 2.5 Implementing the Parser(at home) 30.07.2005 IMPL 4.0 Implementing the Parser(at home) 31.07.2005 DESI 1.0 Reading the SDD for the review on 03.08 01.08.2005 MEET 2.0 Meeting with the project group 01.08.2005 IMPL 2.0 Implementing the Parser(at home) 02.08.2005 IMPL 4.0 Implementing the Parser and fixing bugs 02.08.2005 IMPL 3.0 Implementing the Parser and fixing bugs 03.08.2005 MEET 0.5 Official review of the SDD & meeting 03.08.2005 IMPL 1.5 Fixing bugs and tweaking the parser 03.08.2005 IMPL 1.0 Fixing bugs and tweaking the parser(at home) 04.08.2005 IMPL 3.0 Finishing the Parser and tweaking the prototype 05.08.2005 MEET 1.5 Meeting with the project group 05.08.2005 IMPL 1.5 Adding features to the prototype 05.08.2005 IMPL 3.0 Adding features to the prototype and Parser 08.08.2005 MEET 1.5 Meeting with the project group 08.08.2005 IMPL 1.5 Adding proposal strategy to the parser, repackaging 08.08.2005 IMPL 4.0 Fixing FortranWriter, implementing generateRead-

Data

09.08.2005 MEET 1.0 Meeting with the customer about new example mod-

els

09.08.2005 IMPL 3.0 Writing Input.java 10.08.2005 OTHE 2.0 On Taina’s testing lecture 10.08.2005 IMPL 4.0 Finishing Input, fixing FortranWriter indenting 11.08.2005 MEET 2.0 Meeting with the project group 11.08.2005 IMPL 3.0 Implementing the spatial part of Acceptation.java 12.08.2005 IMPL 2.0 Implementing generateSetFunctional 15.08.2005 MEET 3.0 Official meeting 15.08.2005 TEST 5.0 Writing the Testing document 16.08.2005 TEST 2.5 Meeting & Writing tests for parseVariable 16.08.2005 TEST 5.0 Writing tests for parseVariable and the testing doc 17.08.2005 TEST 2.0 Writing the testing document and discussing features 17.08.2005 TEST 5.0 Ironing out bugs and bettering Output.java 18.08.2005 TEST 2.0 Official meeting & finding a bug 19.08.2005 TEST 2.0 Fixing the bug (functionals now update correctly) 19.08.2005 TEST 3.0 Testrunning on the cancer model, fixing misc. bugs 20.08.2005 OTPR 1.0 Javadocing the classes, writing the manual glossary 22.08.2005 TEST 2.0 Made the MasterTest class and ispelled the manual 23.08.2005 MEET 3.0 Meeting with the project group 23.08.2005 OTPR 4.0 Correcting the manual. 24.08.2005 MEET 3.0 Meeting with the project group and another with the

customer

24.08.2005 TEST 2.0 Writing the testing document 24.08.2005 OTPR 3.0 Making a handy script for project release 25.08.2005 TEST 1.0 Incorporating tests into MasterTest 25.08.2005 OTPR 2.0 Writing the manual and modifying the Generator 25.08.2005 OTPR 4.0 Writing the manual and the cdrom script 26.08.2005 MEET 1.0 Meeting with the project group 26.08.2005 TEST 2.5 Testing the missing values handling in the generated

simulation

26.08.2005 OTPR 1.0 put CD-ROM to the web page and other stuff 26.08.2005 OTPR 2.0 Checked the manual, testing doc, wrote some of the

impl. doc

28.08.2005 TEST 4.0 Fixed some latent bugs and refactored stuff. The bird

model doesn’t work yet.

29.08.2005 MEET 1.0 Meeting with the project group 29.08.2005 OTPR 2.0 Wrote some of the implementation document 30.08.2005 OTPR 2.0 Wrote some of the impl doc 30.08.2005 OTPR 1.0 Reading impl., conc., test. docs for review 31.08.2005 MEET 1.5 Meeting and review of the docs 31.08.2005 OTPR 1.5 Correcting the documents